# Reasoning-Aware Proxy Reward Model using Process Mining

### v2

@@ 4 Experiments
Table 3: Comparison with state-of-the-art RL methods using GSPO + TACReward (3000 training steps)
@@
DeepMath (He et al., 2025b) 83.4 42.6 48.7 34.2 30.0 63.3 17
SkyWork (He et al., 2025a) 87.0 43.4 55.7 46.7 36.7 66.7 12
GRPO-LEAD (Zhang and Zuo, 2025) 84.6 47.4 52.3 40.0 26.7 76.7 13
DRGRPO (Liu et al., 2025c) 80.2 10.3 41.0 40.0 6.7 66.7 21
ExGRPO (Zhan et al., 2025) 82.8 41.0 47.9 23.3 16.7 63.3 13
Eurus-2 (Cui et al., 2025) 79.2 38.6 42.1 26.7 16.7 20.0 2
VeriThinker (Chen et al., 2025b) 80.2 34.9 46.4 30.0 13.3 76.7 14
@@
DRA-GRPO (Chen et al., 2025a) 71.0 26.1 40.4 20.0 23.3 43.3 5
Open-RS3 (Dang and Ngo, 2025) 69.8 26.8 39.7 20.0 20.0 50.0 6
STILL-3 (Min et al., 2024) 72.4 23.2 42.8 20.0 16.7 53.3 3
ExGRPO (Zhan et al., 2025) 71.2 30.9 34.4 10.0 10.0 40.0 10
@@
Figure 3: Reward changes during GSPO + TAC training, smoothed with a 15-step moving average. Left: 7B model. Right: 1.5B model.
Figure 4: Changes in Entropy, Mean Length, TACReward for Fitness vs. Precision vs. Both (240 training steps)
4.1 Experiment Setup
Models & Baselines
We consider two categories of base models. For non-reasoning base models, we used Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct (Qwen et al., 2025), which did not were fine-tuned with RL for reasoning capabilities. For the reasoning-enhanced base models, we use DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), which already possess reasoning capabilities based on distillation.
Benchmarks & Evaluation Metric
We train on the DeepMath-103k (He et al., 2025b) dataset and evaluated seven challenging mathematical reasoning benchmarks: MATH-500 (Hendrycks et al., 2021) MINERVA (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), LiveMathBench (Liu et al., 2025b) and Korean CSAT Math Calculus (KICE, 2025), and AIME 2024–2025 (MAA, n.d.). However, we did not address this problem MATH500 and AIME 2024 as primary evidence because of potential contamination (Wu et al., 2025). Therefore, the results for MATH-500 and AIME 2024 are reported for reference only, and the system prompts that was previously mentioned, was not used. Instead, all of the main quantitative comparisons and conclusions are obtained from the remaining benchmarks. All evaluations used the Pass@1 metric temperature = 0.6, top_p = 0.95, and max_tokens = 16384, following the flashinfer (Ye et al., 2025) framework. The answers are extracted using the DeepMath Evaluation Library  (He et al., 2025b).
RL Settings
All experiments are conducted using the TRL framework (von Werra et al., 2020) on $4\times$ NVIDIA H200 GPUs. We employed the DeepSeek API (Guo et al., 2025) as the reward model that provides both accuracy and conformance reward signals. We used the AdamW optimizer for optimization at a constant learning rate of $1\times 10^{-6}$. The training was performed with a global batch size of 128, utilizing micro-batch sizes of 8 for the 1.5B models and 4 for 7B models, respectively. We set the KL divergence coefficient to zero for unconstrained policy updates. The details of hyperparameters are provided in Appendix A.
4.2 Foundational Policy Gradients Methods with TACReward
Strong Synergy with GSPO.
The most significant improvement was observed when TACReward was combined with GSPO. Although GSPO alone underperforms in this early stage regime, GSPO + TAC increases Avg. * from 17.2 to 32.5 (+89.2%). We attribute this synergy to a granularity match between the optimization objective and reward signal. GSPO defines the importance ratio and updates it at the sequence level, effectively treating each sampled response as a single optimization unit. TACReward aggregates the step-level deviations revealed by the trace alignment into a single sequence-level conformance score that represents the reasoning quality of one response. Conseqently, both components operate under a coherent unit of credit assignment—one reward per response. This allows TACReward to provide more informative and a well-aligned training signal for GSPO and can be exploited more effectively than in token-level or step-mismatched objectives.
General Improvements with RLOO and GRPO.
TACReward provides consistent improvements for RLOO and GRPO, with relative gains of +6.1% and +12.7%, respectively, in Avg. This suggests a broad compatibility with variance-reduced and group-based objectives. Stronger baselines are less sensitive to the additional variance introduced by proxy rewards.
Limited gains with PPO.
Adding TACReward to PPO yielded limited gains and slightly reduced the overall average score (Avg. *) despite improvements in some benchmarks. This behavior can be explained by the inherent characteristics of PPO. These characteristics aim to dense reward settings, and can be adapted to sparse rewards. In this study, PPO was trained under sparse reward regime, in which sensitivity to reward variance and credit assignment becomes more pronounced. Appendix E provides visualizations of reasoning processes with and without TACReward during GSPO training.
4.3 Comparison with State-of-the-Art Models
@@
Based on the controlled comparison in Table 2, we selected GSPO + TACReward as the most effective configuration and optimized it further for evaluation. Subsequently, we compared this configuration with the recent state-of-the-art RL methods. Table 3 compares GSPO + TAC with the recent state-of-the-art RL methods for two distilled backbone models of different sizes. Training was conducted using 3000 training steps. For both model sizes, GSPO + TAC exhibited strong or state-of-the-art performances across the benchmarks. With the 7B backbone, it outperformed all datasets, particularly Olympiad and KSAT 2025, and remained competitive (often better) at 1.5B.
@@
Figure 3 shows that TACReward steadily increases during training, indicating a progressive better conformance between the policy’s reasoning process and teacher’s traces. For both the model sizes, the model learns to produce more accurate answers while more closely aligning the reasoning process with the teacher. Appendix F provides visualizations of reasoning processes as training progresses, indicating that the model’s reasoning steps became more structured and aligned with those of the teacher over time.
Proof of Concept for Using Both Fitness and Precision.
Based on the definitions in (Van Der Aalst, 2016), fitness penalizes the missing essential steps, whereas precision penalizes excessive or unjustified behavior. Through Figure 4, we find that this definition aligns exactly with the expected behavior patterns that we want to encourage during training.
@@
Minerva Olympiad AIME 2025 KSAT 2025 LiveMath Bench
w/ Fit 30.9 30.5 6.7 43.3 10
w/ Prec 35.7 35.7 3.3 46.7 12
w/ Both 37.1 38.2 10.0 53.3 15
First, when only fitness was used, we expected the model to generate as many reasoning steps as possible and to employ a wide variety of activities to avoid penalization for missing essential steps. This is confirmed in Figure 4, where using only fitness (green) results in longer reasoning chains (mean length). However, when only precision is used (purple), the model becomes conservative, generating fewer reasoning steps and using fewer activities to minimize excessive behavior. This is indicated by the shorter mean length and higher entropy. Both components (red) balance these two tendencies, yielding the highest confidence (lowest entropy) and moderate length. Appendix G provides visualizations of reasoning processes, where the process models extracted from the same problem visually demonstrate this assumption. Table 4 further confirms that using either component alone leads to a degraded performance across benchmarks.
5 Discussion
5.1 Design Rationale and Methodological Choices
Role of Mathematical Taxonomy.
@@ 5.1 Design Rationale and Methodological Choices
Minerva Olympiad AIME 2025 KSAT 2025 LiveMath Bench
w/o MRT 37.1 35.6 3.3 43.3 9
w/ MRT 37.1 38.2 10.0 53.3 15
Table 5 indicates that removing MRT degrades performance, with larger drops on multistep benchmarks (e.g., AIME 2025 and KSAT 2025). In this setting, the rewards become less sensitive to structural differences and are dominated by surface variation without an explicit taxonomy. Appendix D provides further discussions on the design rationale behind TACReward.
5.2 Limitations
Computational Overhead.
TACReward includes additional computations for trace formalization, process discovery, and conformance checking.
Table 6: Per-Step Training Time (s) Across Algorithms With/Without TACReward
PPO RLOO GRPO GSPO
w/o TAC 8.0 8.5 8.1 10.9
w/ TAC 39.7 37.9 64.3 67.2
As shown in Table 6, TACReward increases the per-step training time across all algorithms. This overhead is the main trade-off in the proposed approach.
6 Related Works
Recent studies have explored approaches that consider the quality of reasoning processes in a sparse reward policy gradient methods for LRMs (Cui et al., 2025; Fan et al., 2025; Yang et al., 2025; Zhan et al., 2025). However, these methods rely on an indirect estimation or approximation of the overall reasoning quality. For example, metrics such as the final answer correctness, model confidence (e.g., entropy) (Zhan et al., 2025), preferences for optimized-degraded reasoning pairs (Fan et al., 2025), expectations of tree-sampled continuations (Yang et al., 2025), and Process rewards derived from outcome supervision(Cui et al., 2025) were used to estimate the overall reasoning quality. Although these methods capture reasoning quality, it remains challenging to consider the resulting reward signal as reflecting the quality of the individual reasoning steps. In contrast, TACReward is computed based on the step-level deviations in the reasoning process. Consequently, it serves as a proxy for step-level reasoning quality.
@@ 7 Conclusion
Sparse reward policy gradient methods are challenging to incorporate into reasoning tasks, particularly when learning which reasoning steps to perform and in which order. Moreover, providing such annotations is costly, and their quality can significantly affect training outcomes. In this study, we proposed TACReward, a reasoning-aware proxy reward that can be seamlessly integrated into sparse reward policy gradient methods for LRM post-training without additional annotations or architectural changes. Instead of teaching the model’s exact intermediate actions, TACReward measures the structural alignment between the reasoning processes of a policy model and that of a more logically mature teacher. The results showed that accounting for structural similarity improves the sparse reward performance and optimization behaviors induced by fitness and precision match their intended effects. Therefore, TACReward effectively captures key structural properties of reasoning. However, TACReward introduces additional computational overhead. Future studies must focus on reducing API dependence and improving the efficiency.
Impact Statement
This study proposes a TACReward for advancing the field of language model post-training. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Details of Hyperparameters
@@
Appendix B Prompts for Model Training
B.1 Prompt for Generation

### v3

@@ 4 Experiments
In this section, we present the experimental results of integrating TACReward into sparse reward policy gradient methods. The details of experiments are described in Appendix˜B.
4.1 Foundational Policy Gradients Methods with TACReward
@@
Strong Synergy with GSPO.
The most significant improvement was observed when TACReward was combined with GSPO. Although GSPO alone underperforms in this early stage regime, GSPO + TAC increases Avg. * from 17.2 to 32.5 (+89.2%). We attribute this synergy to a granularity match between the optimization objective and reward signal. GSPO defines the importance ratio and updates it at the sequence level, effectively treating each sampled response as a single optimization unit. TACReward aggregates the step-level deviations revealed by the trace alignment into a single sequence-level conformance score that represents the reasoning quality of one response. This allows TACReward to provide more informative and well-aligned training signal for GSPO and can be exploited more effectively than in token-level or step-mismatched objectives.
Compatibility with PPO, RLOO, and GRPO.
TACReward provides consistent improvements for RLOO and GRPO, with relative gains of +6.1% and +12.7%, respectively, in Avg.*, suggesting broad compatibility with variance-reduced and group-based objectives. In contrast, adding TACReward to PPO yields limited gains and slightly reduces the overall average score despite improvements on some benchmarks. We attribute this asymmetry to PPO’s reliance on dense reward signals: under the sparse reward regime considered in this study, PPO becomes more sensitive to reward variance and credit assignment, whereas RLOO and GRPO are inherently designed for variance reduction and thus absorb the additional signal from TACReward more effectively. Appendix˜E provides visualizations of reasoning processes with and without TACReward during GSPO training.
4.2 Comparison with State-of-the-Art Models
Table 3: Comparison with baseline RL methods using GSPO + TACReward (3000 training steps)
@@ 4.2 Comparison with State-of-the-Art Models
DeepMath [12] 83.4 42.6 48.7 34.2 30.0 63.3 17
SkyWork [11] 87.0 43.4 55.7 46.7 36.7 66.7 12
GRPO-LEAD [46] 84.6 47.4 52.3 40.0 26.7 76.7 13
DRGRPO [21] 80.2 10.3 41.0 40.0 6.7 66.7 21
ExGRPO [45] 82.8 41.0 47.9 23.3 16.7 63.3 13
Eurus-2 [6] 79.2 38.6 42.1 26.7 16.7 20.0 2
VeriThinker [5] 80.2 34.9 46.4 30.0 13.3 76.7 14
@@
DRA-GRPO [4] 71.0 26.1 40.4 20.0 23.3 43.3 5
Open-RS3 [7] 69.8 26.8 39.7 20.0 20.0 50.0 6
STILL-3 [24] 72.4 23.2 42.8 20.0 16.7 53.3 3
ExGRPO [45] 71.2 30.9 34.4 10.0 10.0 40.0 10
Refer to captionRefer to caption
@@
Figure 5: Reward dynamics during GSPO + TAC training. Left: Comparison of accuracy reward among GRPO, GSPO, and GSPO + TAC on the 1.5B policy model. Right: TACReward (left axis) and accuracy reward (right axis) over training steps for the 7B (top) and 1.5B (bottom) policy models, smoothed with a 15-step moving average.
@@
Based on the controlled comparison in Table 2, we selected GSPO + TACReward as the most effective configuration and optimized it further for evaluation. Subsequently, we compared this configuration with the recent state-of-the-art RL methods. Table 3 compares GSPO + TAC with the recent state-of-the-art RL methods for two distilled backbone models of different sizes. Training was conducted for 3000 training steps. For both model sizes, GSPO + TAC exhibited strong or state-of-the-art performance across the benchmarks. With the 7B backbone, it outperformed all baselines, particularly Olympiad and KSAT 2025, and remained competitive (often better) at 1.5B.
@@
The comparison of reward dynamics among GRPO, GSPO, and GSPO + TAC in the 1.5B policy model in Figure˜5 (Left) illustrates that GSPO + TAC achieves stable and high reward trajectories throughout training. Moreover, TACReward steadily increases during training, indicating a progressive better conformance between the policy’s reasoning process and the teacher’s traces. For both model sizes, the model learns to produce more accurate answers while more closely aligning the reasoning process with the teacher (Right). Appendix˜F provides visualizations of reasoning processes as training progresses, indicating that the model’s reasoning steps became more structured and aligned with those of the teacher over time.
4.3 Analysis
Table 4: Performance comparison of distilled models against the DeepSeek R1 teacher across reasoning benchmarks. Relative degradation from the teacher is shown in parentheses.
@@ 4.3 Analysis
Teacher Size Teacher Acc. MATH 500 MINERVA Olympiad AIME 2024 AIME 2025 KSAT 2025 LiveMath Bench
DeepSeek-R1-Distill-Qwen-1.5B
DeepSeek R1 685B 94.17% 49.8 19.7 20.6 0 0 30 12
R1-Distill-LLaMA 70B 86% ($8.7\%$) 45.8 19.1 18.4 0 0 30 10
R1-Distill-Qwen-2.5 32B 68% ($21.8\%$) 44.8 17.6 17.3 0 0 26.7 12
R1-Distill-Qwen-2.5 14B 67.8% ($28.0\%$) 44.2 15.4 17.3 3.3 0 23.3 12
Sensitivity to Teacher Quality.
A natural concern is whether TACReward depends on access to a frontier-scale teacher. Table˜4 compares the performance by replacing DeepSeek-R1 (671B) with progressively smaller and less accurate teachers, while keeping all other components fixed. Despite teacher accuracy on DeepMath-103k dropping by up to 28%, the resulting policy’s performance degrades only marginally on most benchmarks. This indicates that TACReward does not require the teacher to be correct on every problem; what matters is the structural pattern of the reasoning trace, which remains informative even when the teacher’s final answer is wrong.
Table 5: Policy performance when varying the trace extractor model.
@@
Extractor MATH 500 MINERVA Olympiad AIME 2024 AIME 2025 KSAT 2025 LiveMath Bench
DeepSeek-V3.2 49.8 19.7 20.6 0.0 0.0 30.0 12
Gemini-2.5-Flash 49.8 19.7 20.2 3.3 0.0 26.7 12
GPT-5.4-nano 50.0 20.3 20.8 3.3 0.0 30.0 12
Sensitivity to Trace Extractor.
In TACReward, a general-purpose model is utilized to convert raw reasoning responses into activity sequences (Section˜3). To assess whether this step is bottlenecked by extractor capacity, we replace DeepSeek-V3.2 [18] with two alternatives spanning different model families: Gemini-2.5-Flash and GPT-5.4-nano. The latter is the smallest member of the GPT-5.4 family. As shown in Table 5, all three extractors yield nearly identical policy performance, with GPT-5.4-nano matching or slightly exceeding our default on every benchmark. Trace formalization is thus robust to the choice of extractor as long as the model can follow the 20-activity taxonomy.
Proof of Concept for Using Process Mining.
Based on the definitions in [38], fitness penalizes the missing essential steps, whereas precision penalizes excessive or unjustified behavior. Through Figure˜6, we found that this definition aligns exactly with the expected behavior patterns that we want to encourage during training.
Figure 6: Changes in Entropy, Mean Length, and TACReward for Fitness vs. Precision vs. Both (240 training steps)
When only fitness was used, we expected the model to generate as many reasoning steps as possible and to employ a wide variety of activities to avoid penalization for missing essential steps. This is confirmed in Figure˜6, where using only fitness (green) results in longer reasoning chains (mean length). However, when only precision is used (purple), the model becomes conservative, generating fewer reasoning steps and using fewer activities to minimize excessive behavior. This is indicated by the shorter mean length and higher entropy. Both components (red) balance these two tendencies, yielding the highest confidence (lowest entropy) and moderate length. Appendix˜G visualizes this behavior, and Table˜7 confirms that using either component alone degrades performance across benchmarks.
Limitations
TACReward requires a task-specific taxonomy: the current 20-activity scheme is tailored to mathematical problem solving, and other reasoning domains may demand a new taxonomy together. We further note that our experiments use high-cost API-based teacher and extractor models to showcase the full potential. In practice, one can use smaller models for both roles, and the sensitivity analyses suggest that this would not significantly degrade performance. However, the exact performance trade-offs and optimal configurations when using smaller models remain to be explored.
For additional analyses, Appendix˜C provides further results and discussions.
@@ 5 Conclusion
Sparse reward policy gradient methods are challenging to incorporate into reasoning tasks, particularly when learning which reasoning steps to perform and in which order. Moreover, providing such annotations is costly, and their quality can significantly affect training outcomes. In this study, we proposed TACReward, a reasoning-aware reward model that can be seamlessly integrated into sparse reward policy gradient methods for LRM post-training without additional annotations or architectural changes. Instead of teaching the exact intermediate actions, TACReward measures the structural alignment between the reasoning processes of a policy model and that of a more logically mature teacher. The results showed that accounting for structural similarity improves the sparse reward performance and optimization behaviors induced by fitness and precision match their intended effects.
For future works, the underlying mechanism of TACReward is general and we expect it to apply to a broader range of reasoning domains such as code generation. We argue that a particularly promising direction is to extend TACReward to domains where the underlying process is well-defined. In such settings, a teacher model is not strictly necessary: the reference trace can be specified directly from domain knowledge, and TACReward can serve as a structural guardrail that evaluates whether the policy’s reasoning conforms to the prescribed procedure. For instance, in loan assessment, the reference trace can encode the standard sequence of steps a credit analyst is expected to perform, and TACReward can guide the policy to internalize a reasoning process appropriate for the task.
Appendix A Integrating TACReward with Sparse Reward Policy Gradients Methods
In this section, we demonstrate how TACReward can be easily integrated into the foundational sparse reward policy gradient methods: PPO, RLOO, GRPO, and GSPO.
Notations
TACReward is denoted as $\left\{r^{TAC}(x,y_{i})\right\}_{i=1}^{G}$. We set $\left\{r^{acc}(x,y_{i})\right\}_{i=1}^{G}$ as the accuracy reward and $\left\{r^{fmt}(x,y_{i})\right\}_{i=1}^{G}$ as follows: thinking format reward for each response $\left\{y_{i}\right\}_{i=1}^{G}$. The total reward for each query-response pair is denoted as $\left\{R(x,y_{i})\right\}_{i=1}^{G}$ as described in Section˜2. In cases where multiple responses are not generated for the same query $G=1$ and are simply denoted as $R(x,y)$.
A.1 Proximal Policy Optimization (PPO) with TACReward
PPO [30] generates a single response for each input query (i.e., $G=1$) and update the policy model based on the rewards obtained from the response. For an input $x\sim\mathcal{D}$ and a sampled response $y=\left\langle y_{1},y_{2},\dots,y_{\left\lvert y\right\rvert}\right\rangle\sim\pi_{\theta_{\mathrm{old}}}(\cdot|x)$, TACReward can be incorporated into PPO by simply adding it as follows:
A.2 REINFORCE with Leave-One-Out (RLOO) with TACReward
RLOO [1] generates a group of $G$ responses for each input query (i.e., $G>1$) and updates the policy using a leave-one-out baseline. Given an input $x\sim\mathcal{D}$ and set of sampled responses $\{y_{i}=\langle y_{i,1},\dots,y_{i,|y_{i}|}\rangle\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|x)$, TACReward can be integrated into RLOO by adding it to the reward for each response as follows:
Using the score function estimator, the RLOO policy gradient is
A.3 Group-based Policy Optimization with TACReward
GRPO and GSPO bypass the need for the value model by computing the group-relative advantage of each response within a response group for the same query. Partially borrowed from [19], the group-relative advantages for the responses $\left\{y\right\}^{G}_{i=1}$ are computed as:
where $\epsilon$ is a hyperparameter for the clipping range, and the importance ratio $\omega_{i,t}(\theta)$ is $\frac{\pi_{\theta}\left(y_{i,t}|x,y_{i,<t}\right)}{\pi_{\theta_{old}}\left(y_{i,t}|x,y_{i,<t}\right)}$ The GSPO simplifies this objective as follows:
Appendix B Details of Experiments
B.1 Hyperparameters Setting
@@
B.2 Experimental Settings
Models & Baselines
We consider two categories of base models. For non-reasoning base models, we used Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [28], which did not fine-tuned with RL for reasoning capabilities. For the reasoning-enhanced base models, we use DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B [9], which already possess reasoning capabilities based on distillation.
Benchmarks & Evaluation Metric
We train on the DeepMath-103k [12] dataset and evaluated seven challenging mathematical reasoning benchmarks: MATH-500 [13] MINERVA [16], OlympiadBench [10], LiveMathBench [20] and Korean CSAT Math Calculus [14], and AIME 2024–2025 [22]. However, we did not address this problem MATH500 and AIME 2024 as primary evidence because of potential contamination [42]. Therefore, the results for MATH-500 and AIME 2024 are reported for reference only, and the system prompts that was previously mentioned, was not used. Instead, all of the main quantitative comparisons and conclusions are obtained from the remaining benchmarks. All evaluations used the Pass@1 metric temperature = 0.6, top_p = 0.95, and max_tokens = 16384, following the flashinfer [44] framework. The answers are extracted using the DeepMath Evaluation Library  [12].
RL Settings
All experiments are conducted using the TRL framework [39] on $4\times$ NVIDIA H200 GPUs. We employed the DeepSeek API [9] as the reward model that provides both accuracy and conformance reward signals. We used the AdamW optimizer for optimization at a constant learning rate of $1\times 10^{-6}$. The training was performed with a global batch size of 128, utilizing micro-batch sizes of 8 for the 1.5B models and 4 for 7B models, respectively. We set the KL divergence coefficient to zero for unconstrained policy updates. The details of hyperparameters are provided in Appendix˜B.
Appendix C Additional Analysis
C.1 Computational Complexity of the optimal alignment.
We analyze the computational cost of TACReward as the length of the reasoning increases. Let $n=|\sigma^{\text{ref}}|$ denote the length of the reference trace, $m=|\sigma_{i}|$ the length of the $i$-th policy trace, and $|\Sigma|$ the size of the activity set. In our setting, $|\Sigma|\leq 20$ (Section 3).
Optimal alignment.
Computing the optimal alignment $\gamma^{*}_{i}$ between a reference trace $\sigma^{\text{ref}}$ and a policy process model $\mathcal{M}_{i}$ can be cast as an $A^{*}$ search over the synchronous product of the two [37]. In the general case this search is exponential in the trace length, but three properties of our setting reduce it substantially: (i) the activity taxonomy is fixed, so $|\Sigma|$ is a constant; (ii) the number of reachable intermediate states in each process model is bounded by a constant, since each $\mathcal{M}_{i}$ is discovered from a single trace; and (iii) the search graph contains no branching beyond local model moves. Under these conditions, the cost of finding $\gamma^{*}_{i}$ reduces to $O(n\log n)$.
Output reward.
Once $\gamma^{*}_{i}$ is obtained, the fitness score requires summing local move costs along the alignment, which takes $O(n+m)$. The precision score iterates over events in the reference trace and inspects the set of enabled activities at each step, costing $O(n\cdot|\Sigma|)$. Since $|\Sigma|$ is constant, both scores reduce to $O(n)$, and the harmonic mean in Equation 14 adds only constant overhead. The overall cost of computing $r_{i}^{TAC}$ is therefore dominated by the alignment step at $O(n\log n)$.
C.2 TAC as a Proxy for Step-level Reward
The alignment $\gamma_{i}^{*}$ decomposes the discrepancy between the policy and teacher reasoning traces into individual moves, each incurring local cost $\delta_{m}(\cdot)$. The total deviation cost performs implicit credit assignment at the reasoning transition level, even though the final reward is scalar. A log-only move $(a_{ref},\gg)$ indicates that the policy model omits the reasoning activity that is present in the teacher’s trace and corresponds to a missing logical step. Conversely, the model-only move $(\gg,a_{m})$ reflects redundant or irrelevant reasoning activities that are not supported by the teacher’s process.
C.3 Why Single-Trace Can Be Effective.
C.4 Why Response-Level Comparison Is Insufficient.
A natural question is whether trace formalization can be skipped, and whether reasoning responses can be directly compared as sequences. Given two token sequences, $y$ and $y^{\prime}$, the classical edit distance [40] is
that measures surface-level similarity via insertions, deletions, and substitutions. However, $d_{\mathrm{ed}}$ measures only surface edit operations on raw text, and cannot reliably capture the structural properties of reasoning (e.g., missing or spurious transitions) that are distinct from benign re-orderings [32]. Trace formalization separates reasoning structure from linguistic realization, enabling a meaningful structural comparison.
C.5 On Imperfect Reference Traces.
In this study, our approach does not assume that the reference model is correct. This assumption is shared using several existing post training methods, including GRPO and GSPO. These methods rely on reward models or KL-divergence regularization with respect to a reference policy, despite these references being imperfect [25]. The reference model must be more mature than the policy being optimized, providing a relatively stable and coherent reasoning process. Under this assumption, conformance-based rewards suppress early stage randomness and structurally invalid reasoning, whereas outcome-based rewards continue to guide convergence as the training progresses.
C.6 Role of Reward Components.
TACReward consists of two components: fitness and precision. To investigate the contribution of each component, we conducted an ablation study by training the policy model with only fitness, only precision, and both components. The results are summarized in Table 7.
Minerva Olympiad AIME 2025 KSAT 2025 LiveMath Bench
w/ Fit 30.9 30.5 6.7 43.3 10
w/ Prec 35.7 35.7 3.3 46.7 12
w/ Both 37.1 38.2 10.0 53.3 15
C.7 Role of Mathematical Taxonomy.
Minerva Olympiad AIME 2025 KSAT 2025 LiveMath Bench
w/o MRT 37.1 35.6 3.3 43.3 9
w/ MRT 37.1 38.2 10.0 53.3 15
Table 8 indicates that removing MRT degrades performance, with larger drops on multistep benchmarks (e.g., AIME 2025 and KSAT 2025). In this setting, the rewards become less sensitive to structural differences and are dominated by surface variation without an explicit taxonomy.
Appendix D Prompts for Model Training
D.1 Prompt for Generation